40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 776     Entry time: Thu Jul 31 11:19:30 2008     In reply to: 748     Reply to this: 782
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ Resonance search trial 
Last night, I tried to find the resonance of Yarm by sweeping the frequency of the injection beam.

A strong beat was present at LT_NPRO=48.7856[C_deg], the power coupling of the injection beam was estimated to be 35%. 
(Vmax_beat = 1.060[V], Vmin_beat = 0.460[V], Vno_inject = 0.664[V])

The Yarm was locked and the alignment script was executed. The PLL between the PSL beam and the injection beam was 
somehow locked.

I tried to scan the freq offset (f_PLL) at around 3.88MHz first, then at around 15.52MHz. They are supporsed to be the 
first and fourth FSR of the Yarm cavity. The Yarm transmitted power (DC) was observed to find the resonance of the 
injection beam. It would have been better to use the RF power, but so far I didnot have the RF PD prepared at the end 
transmission. I just used the DC power.

I think I saw the increase of the transmitted power by 10%, at f_PLL = 15.517 +/- 0.003 [MHz]. This corresponds to the 
arm cavity length of 38.640 +/- 0.007 [m]. The previous measurement was not so bad!

Y-arm length
e-log             length [m]
-----------------------
556(2008-Jun-24)  38.70    +/- 0.08    Cavity swinging measurement
556(2008-Jun-24)  38.67    +/- 0.03    tape & photo
This              38.640   +/- 0.007

However, I had difficulties to have more precise measurement mainly because of two reasons:
o The PLL servo is too naive, and the freqency stability of the inj beam is not enough.
  The injected beam should have the linewidth (=freq stability) narrower than the cavity linewidth.

o The PLL servo may experience change of the transfer function at around the resonance. The PLL works the other 
frequencies. However, close to the resonance, it starts to be unstable.

So the next stuffs we should do is 
o Build the PLL just using the incident beams to the ifo, not by the reflected beams.
o Build sophisticated servo to have better frequency stability.
o RF PD at the transmission.

Left the lab with Yarm locked, flipper down, shutter for the NPRO closed.
ELOG V3.1.3-